# Turnip: A “Nondeterministic” GPU Runtime with CPU RAM Offload

### v1

@@ Abstract
An obvious way to alleviate memory difficulties in GPU-based AI computing is via CPU offload, where data are moved between GPU and CPU RAM. While CPU offload is useful, it can greatly slow down a computation due to the relatively transfer rate between CPU RAM and GPU RAM. To address this, overlapping memory transfer and compute is a necessity, but this asynchronicity introduces nondeterminacy, and hence it is impossible to know beforehand what is the best order of operations. We describe Turnip, which is a system for running AI computations using CPU offload, designed to handle this nondeterminacy. The key innovation in Turnip is the compilation of the AI computation into a dependency graph that gives the Turnip runtime freedom to run operations such as GPU kernel calls in many different orders; at runtime, Turnip chooses the best order on-the-fly, in response to real-time events. We find that Turnip outperforms standard, PyTorch-based systems supporting constrained GPU RAM by a significant margin, and also avoids out-of-memory errors in a severely memory constrained environment.
@@ 1 Introduction
In this paper, we propose Turnip (short for “nondeTerministic gpU RuNtime wIth cPu offload”) which is a runtime for multi-GPU servers, designed to systematically support CPU RAM offload. The key innovation of Turnip is its combination of a pre-computed memory access plan called a memgraph with a “nondeterministic,” event-driven system runtime. A memgraph is a dependency graph where vertices represent tasks (such as the execution of a GPU kernel to perform a small part of of attention computation in a layer of a large language model) and edges represent data or memory dependencies. Any execution order that respects the dependencies in the memgraph is valid, and tasks are dispatched at any time that their dependencies have been met and the appropriate resources are free. Thus, two executions of the same memgraph may lead to different sequences of operations being executed on a GPU, or different sequences of tensors being paged to CPU RAM—hence the non-determinacy. However, the dependencies in the memgraph are such that the final output is always correct, no matter the execution order. Turnip’s event-driven, fully asynchronous runtime is unique. Because operations can be dispatched whenever the dependencies are fulfilled and are not constrained to any specific ordering, it lowers the chance that any GPU will be stalled waiting for a memory transfer to complete. If one task cannot run due to an un-met dependency in the memgraph, it is possible that there is another task that can run.
@@ 3 Related Work
The more targeted approach is taken by the DeepSpeed project (Deepspeed, ) and the various ZeRO optimizations. For transformers and other, similar models, DeepSpeed inference (which includes ZeRO-Inference) (Aminabadi et al., 2022) has two key ideas. First, DeepSpeed inference “offload[s] some activation from GPU to CPU memory while not in use.” Second, DeepSpeed inference “pins the model weights either in DRAM (if large enough) or NVMe, and streams each layer into GPU memory for computation when needed.” FlexGen (Sheng et al., 2023) seeks to use a variety of methods to speed transformer inference given limited hardware, including model weight offload to CPU, quantization (Yao et al., 2022; Frantar et al., 2022), and sparse attention (Child et al., 2019). The latter two ideas are orthogonal to the ideas in this paper. For CPU offload, FlexGen optimizes a “zig-zag” block scheduling that works through a tansformer layers and sequences in the batch, offloading and reloading the KV-cache (Pope et al., 2023) and model weights. PagedAttention (Kwon et al., 2023) deals with low memory utilization in transformers, developing a paging system for the KV-cache.
@@
Given such a taskgraph, Turnip first compiles the taskgraph into a memgraph, which it will eventually execute. Like a taskgraph, a memgraph is also a directed acyclic graph. Every vertex in the original taskgraph will be present in a corresponding memgraph. Further, the compilation process may add additional offload and reload operations that move memory from GPU RAM to CPU RAM, and vice versa. During the compilation process, the output associated with every vertex in the memgraph is mapped to a memory location. Unlike the input taskgraph, the memgraph is not a dataflow graph; it is a dependency graph. If there is an edge from $v_{1}$ to $v_{2}$, it means that $v_{2}$ depends on $v_{1}$ and $v_{2}$ may not execute until after $v_{1}$ has executed. In a memgraph, there are two types of dependencies. One is a data dependency, which is inherited from the taskgraph (or is created via the addition of an offload or reload; see below). The second is a memory dependency, which is added to ensure that there are no race conditions in the graph. A race condition occurs when there is some vertex for which two valid executions of the graph may produce a different output. This can happen when two vertices write to the same memory location, and it is possible for a third vertex to read either outupt, depending upon the execution order.
@@ 8 Experiments
Figure 10: Time taken for LLaMA first-token (prefill) inference. “OOM” is out-of-memory.
Figure 11: Training time per batch for LoRA training, 7B param model, 4$\times$×P100 GPU server.
@@
Turnip is implemented in C++, with most GPU kernels generated using Nvidia cuTensor. Experiments were conducted on two machines: (i) an older 4 $\times$ P100 GPU server (16 GB RAM each) and 22, 64GB DDR4 2666MHz CPU RAM modules, for a total of 1.3TB of RAM, and (ii) an Amazon Web Services p4d.24xlarge instance, equipped with eight A100 GPUs (40 GB RAM each) and 1.15TB of RAM. We were particularly interested in seeing the ability of Turnip to operate in a difficult environment with extremely limited GPU RAM, hence the P100 GPUs, with only 64 GB of GPU RAM total on the server. A key question is: can software help bridge the gap—particularly the lack of GPU RAM—between older and newer hardware? The tasks evaluated are:
@@
(1) LLaMA first token inference. Our experiments target “first token” inference (also known as ”prefill”): How long does it take to produce the first output token, given an input prompt? Prefill is exceedingly expensive in terms of the memory required, scaling quadratically with the size of the prompt. On both machines, we run Turnip, ZeRO Inference (Aminabadi et al., 2022) (using weight partitioning and model weight offload), and FlexGen (Sheng et al., 2023). Note that these are PyTorch-based systems, whereas Turnip is not. For FlexGen, we use two settings: both use 50/50 CPU/GPU RAM for the model weight cache; one uses full CPU offload for activations (“Act. off.”), and one does not. As an ablation, we also run Turnip (“Fixed execution”), constrained so that it executes vertices in the order of the simulation used to build the memgraph (thus nondeterminacy is mostly removed from the runtime; see the discussion below). For each option, we test the following prefill input sequence lengths: 1K, 2K, 4K, 8K, 16K, and 32K tokens. All model weights and computations were performed at half precision (though FlexGen uses even lower precision for some computations). A subset of the results are given in Figure 10 (the rest are in the Appendix). For each experiment, we give the time required to produce the first token. “OOM” inidcates running out-of-memory.
@@
(2) LoRA training for LLaMA. We also experiment with LoRA training (Hu et al., 2021). We use a LoRA rank of 16, and train LoRA adaptors for the $K$, $V$, $Q$, and feedforward mapping matrices. Here we run Turnip and ZeRO Infinity (Rajbhandari et al., 2021); ZeRO is executed using all three “stages” (gradient partitioning, model weight partitioning, optimizer state partitioning) as well as CPU offload. Both Turnip and ZeRO use checkpointing during the forward pass to reduce the memory footprint. We again run the Turnip ablation, where execution order is constrained. We measure the time it takes to run the forward and backward pass for one batch, with varying batch sizes (1, 4, 16, 64) and sequence lengths (1K, 2K, 4K, and 8K). All model weights are stored as single precision (32 bits). A subset of the results are given in Figure 11 (other results are in the appendix).
Discussion. Turnip performs significantly better than ZeRO Inference, FlexGen, and ZeRO Infinity, both in terms of runtime, and the ability to handle the severely constrained GPU memory. Among different workloads, Turnip is always the fastest, and avoids most OOM errors. In the most memory-constrained inference experiments (P100 GPUs, 65B parameter model) Turnip was the only method able to go beyond a 2K sequence length without an OOM error. Further, Turnip was able to train the 65B parameter model in many cases on the P100 server, with only 64GB of GPU RAM (see Figure 13 in the Appendix); ZeRO Infinity could handle only a single 1K sequence at a time. One key reason for the robustness is the way in which the sub-computations making up the same, larger operation are decoupled and executed in Turnip. If multiple small matrix multiplies are used to implement a single large matrix multiply (as in Figure 3) they need not run together, en-masse. In a severely memory-constrained environment, Turnip can force them to be run in sequence and offloaded, via the introduction of memory dependencies. The other systems do not do this, and perform layerwise execution of the transformer. This is why FlexGen with CPU offload does not help. As we are performing inference on a single sequence, offloading an activation is not useful (FlexGen’s offload targets CPU offload for the different sequences in a batch, and cannot partially offload the activations for a single sequence).
A secondary finding is that forcing a particular execution order in Turnip (“Fixed execution”) resulted in up to a 3$\times$ slowdown compared to the nondeterministic order. Note that even when nondeterminstic execution is “turned off,” Turnip is still much faster than the PyTorch-based systems, even with ample GPU RAM. One reason for this may be that due to the architecture of Turnip, it is very difficult to completely “turn off” nondeterminism during aggregation—Turnip does not use a synchronous all-reduce, for example, as the other systems do, and instead uses a series of asynchronous, partial summations. This remaining nondeterminism may help explain the still-significant performance advantage of the “Fixed execution” option. Additional experimental results and discussion are given in the Appendix.
@@ 9 Conclusions
Appendix A Additional Experimental Results and Discussion
Figure 12: Time taken for LLaMA first-token (prefill) inference. 7B parameters, 4 $\times$ P100. “OOM” is out-of-memory.
Figure 13: Training time per batch for LoRA training, 65B param model, 4 $\times$ P100 GPU server.
s Refer to caption
Figure 14: Training time per batch for LoRA training, 7B param model, 8 $\times$ A100 GPU server.
Figure 15: Training time per batch for LoRA training, 65B param model, 8 $\times$ A100 GPU server.
Here we give some additional experimental results not found in the main body of the paper, as well as some additional discussion. Figure 12 gives inference results for the P100 server (7B params). Figure 13 gives training results for the 65B parameter model, on the P100 server. Figure 14 gives results for training the 7B parameter model on the A100 server. Figure 15 gives results for training the 65B parameter model on the A100 server.
On training, we were initially surprised to find that there is a relatively small difference in performance between the 4 $\times$ P100 and the 8 $\times$ A100 servers. For example, for batch size 64, 1K sequence length, 7B parameter model, the P100 server took 174 seconds per batch (for Turnip) and 35 seconds per batch on the A100 server. That is, the P100 server is only around 5×5$\times$5 × slower, despite the fact that it has half the number of GPUs, and the GPUs are two generations older. However, upon reflection, this is not that surprising. Much of the increase in compute power in the never GPUs is related to increased 16-bit performance, which is not particularly useful for training. One immediate effect of this is that it magnifies the importance of software, as the performance of the servers for training is not that different—the real problem is the lack of GPU RAM on the P100 server—and a software that can deal with this lack of GPU RAM and cut out as much inefficiency as possible can bridge the gap between the two hardwares. For example, note that Turnip processes a batch of 16, 1K sequences on the P100 server (7B parameter model) in 43 seconds. ZeRO Infinity processes the same 16, 1K sequences on the A100 server in 48 seconds—ZeRO Infinity is actually five seconds slower per batch on the newer, A100 server, compared to Turnip on the P100 server.
Note that on the inference experiments, FlexGen has a relatively flat curve as the sequence length increases. It starts out slower than ZeRO at small sequence lengths, but does not slow down as much at longer sequence lengths. One reason for this may be the various methods embedded in FlexGen (such as a faster attention computation) for speeding up the computation. At smaller sequence lengths, the overhead is high, but the overhead may become less of a factor at larger sequences. Still, Turnip is always faster than both ZeRO and FlexGen—though to be fair, FlexGen specifically targets batch processing, and all of our experiments are single-sequence. Also note that for all experiments, Turnip implements a “vanilla” computation—we use the LLaMA attention mechanism as implemented by LLaMA, and use 32-bit computation everywhere for training, and 16-bit computation everywhere for inference. No techniques such as quantization are used to save space and/or speed inference.
Appendix B Turnip Engine Details
@@
Appendix C Ensuring the “Unobtrusiveness” of Dependencies
Even under nondeterministic execution, memory dependencies can cause periods of time where a GPU is not used, as the GPU is blocked on a memory transfer or on a dependency related to two tensors being mapped to the same location in GPU RAM. The likelihood that dependencies added to the memgraph will cause such stalls can be reduced by a careful implementation of the various malloc variants. For example, when simMallocOffld searches for a victim, we search for the victim whose next use is furthest in the future—that is, closest to the end of $V$ (in the general case where tensors have different sizes and there may be more than one victim, we seek to maximize the minimum age of any evicted tensor). When simMalloc finds an open slot for a tensor, it should choose the available slot whose last use was furthest in the past (closest to the beginning of $V$).

### v3

@@ Abstract
An obvious way to alleviate memory difficulties in GPU-based AI computing is via CPU offload, where data are moved between GPU and CPU RAM. While CPU offload is useful, it can greatly slow down a computation due to the relatively slow transfer rate between CPU RAM and GPU RAM. To address this, overlapping memory transfer and compute is a necessity, but this asynchronicity introduces nondeterminacy, and hence it is impossible to know beforehand what is the best order of operations. We describe Turnip, which is a system for running AI computations using CPU offload, designed to handle this nondeterminacy. The key innovation in Turnip is the compilation of the AI computation into a dependency graph that gives the Turnip runtime freedom to run operations such as GPU kernel calls in many different orders; at runtime, Turnip chooses the best order on-the-fly, in response to real-time events. We find that Turnip outperforms standard, PyTorch-based systems supporting constrained GPU RAM by a significant margin, and also avoids out-of-memory errors in a severely memory constrained environment.
@@ 1 Introduction
In this paper, we propose Turnip (short for “nondeTerministic gpU RuNtime wIth cPu offload”) which is a runtime for multi-GPU servers, designed to systematically support CPU RAM offload. The key innovation of Turnip is its combination of a pre-computed memory access plan called a memgraph with a “nondeterministic,” event-driven system runtime. A memgraph is a dependency graph where vertices represent tasks (such as the execution of a GPU kernel to perform a small part of attention computation in a layer of a large language model) and edges represent data or memory dependencies. Any execution order that respects the dependencies in the memgraph is valid, and tasks are dispatched at any time that their dependencies have been met and the appropriate resources are free. Thus, two executions of the same memgraph may lead to different sequences of operations being executed on a GPU, or different sequences of tensors being paged to CPU RAM—hence the non-determinacy. However, the dependencies in the memgraph are such that the final output is always correct, no matter the execution order. Turnip’s event-driven, fully asynchronous runtime is unique. Because operations can be dispatched whenever the dependencies are fulfilled and are not constrained to any specific ordering, it lowers the chance that any GPU will be stalled waiting for a memory transfer to complete. If one task cannot run due to an un-met dependency in the memgraph, it is possible that there is another task that can run.
@@ 3 Related Work
The more targeted approach is taken by the DeepSpeed project (Deepspeed, ) and the various ZeRO optimizations. For transformers and other, similar models, DeepSpeed inference (which includes ZeRO-Inference) (Aminabadi et al., 2022) has two key ideas. First, DeepSpeed inference “offload[s] some activation from GPU to CPU memory while not in use.” Second, DeepSpeed inference “pins the model weights either in DRAM (if large enough) or NVMe, and streams each layer into GPU memory for computation when needed.” FlexGen (Sheng et al., 2023) seeks to use a variety of methods to speed transformer inference given limited hardware, including model weight offload to CPU, quantization (Yao et al., 2022; Frantar et al., 2022), and sparse attention (Child et al., 2019). The latter two ideas are orthogonal to the ideas in this paper. For CPU offload, FlexGen optimizes a “zig-zag” block scheduling that works through a transformer layers and sequences in the batch, offloading and reloading the KV-cache (Pope et al., 2023) and model weights. PagedAttention (Kwon et al., 2023) deals with low memory utilization in transformers, developing a paging system for the KV-cache.
@@
Given such a taskgraph, Turnip first compiles the taskgraph into a memgraph, which it will eventually execute. Like a taskgraph, a memgraph is also a directed acyclic graph. Every vertex in the original taskgraph will be present in a corresponding memgraph. Further, the compilation process may add additional offload and reload operations that move memory from GPU RAM to CPU RAM, and vice versa. During the compilation process, the output associated with every vertex in the memgraph is mapped to a memory location. Unlike the input taskgraph, the memgraph is not a dataflow graph; it is a dependency graph. If there is an edge from $v_{1}$ to $v_{2}$, it means that $v_{2}$ depends on $v_{1}$ and $v_{2}$ may not execute until after $v_{1}$ has been executed. In a memgraph, there are two types of dependencies. One is a data dependency, which is inherited from the taskgraph (or is created via the addition of an offload or reload; see below). The second is a memory dependency, which is added to ensure that there are no race conditions in the graph. A race condition occurs when there is some vertex for which two valid executions of the graph may produce a different output. This can happen when two vertices write to the same memory location, and it is possible for a third vertex to read either output, depending upon the execution order.
@@ 8 Experiments
Figure 10: Time for LLaMA first token (prefill) inference, A100 server. “OOM” is out-of-memory.
Figure 11: Time for LLaMA first token (prefill) inference, P100 server. “OOM” is out-of-memory.
@@
The system is implemented in C++, with most GPU kernels generated using Nvidia cuTensor. Experiments were conducted on two machines: (i) an older 4 $\times$ P100 GPU server (16 GB RAM each) and 22, 64GB DDR4 2666MHz CPU RAM modules, for a total of 1.3TB of RAM, and (ii) an Amazon Web Services p4d.24xlarge instance, equipped with eight A100 GPUs (40 GB RAM each) and 1.15TB of RAM. We were particularly interested in seeing the ability of Turnip to operate in a difficult environment with extremely limited GPU RAM, hence the P100 GPUs, with only 64 GB of GPU RAM total on the server. Key quesitons are: Can software help bridge the gap—particularly the lack of GPU RAM—between older and newer hardware? Can Turnip facilitate model training and inference in a situation with limited RAM?
Figure 13: Comparing Turnip and ZeRO Infinity for LoRA training.
@@
(1) LLaMA first token inference. Our experiments target “first token” inference (also known as ”prefill”): How long does it take to produce the first output token, given an input prompt? We focus on prefill as it is exceedingly expensive in terms of the memory required, scaling quadratically with the size of the prompt. On both machines, we run Turnip, ZeRO Inference (Aminabadi et al., 2022) (using weight partitioning and model weight offload), and FlexGen (Sheng et al., 2023). Note that these are PyTorch-based systems, whereas Turnip is not. For FlexGen, we use full CPU offload for activations. All testing is done using batched input, as batching is required for FlexGen and ZeRO (as Turnip simply runs a dataflow graph, it is agnostic to batching). For the smallest batch sizes considered, we test prefill input sequence lengths: 1K, 2K, 4K, 8K, and 16K tokens. For larger batches we use 1K, 2K, 4K and 8K. For Turnip, all model weights and computations were performed using 16-bit floating points, though FlexGen uses very low precision arithmetic to save RAM and speed compute. Results for the A100 GPU server are given in Figure 10. Results for the P100 GPU server are Given in Figure 11.
One of the advantages of Turnip is that it executes arbitrary dataflow graphs in limited memory. Thus, as long as a computation is appropriately decomposed to run on multiple GPUs, Turnip can execute it. This means, for example, that Turnip does not need to perform inference over batches of input sequences and supports arbitrary combinations of model and data parallelism (unlike FlexGen and in ZeRO Inference). While batching tends to increase computational efficiency, the RAM used by a large batch means it is not possible to run inference over long sequences in limited memory (batching precludes that all 320GB of GPU RAM on be dedicated to prefill for a single long sequence). To investigate the ability of Turnip to perform inference over a single long sequence, we test sequences sizes of up to 32K tokens, on both GPU servers and on both the 7B and 65B parameter models. Results are shown in Figure 12.
Experiment Avg. Speedup
A100 Inference 4.02%
P100 Inference 6.45%
A100 Training 13.4%
P100 Training 14.5%
Figure 14: Observed speedup due to nondeterministic ordering, compared to (partially) deterministic ordering.
@@
(2) LoRA training for LLaMA. We also experiment with LoRA training (Hu et al., 2021). We use a LoRA rank of 16, and train LoRA adaptors for the $K$, $V$, $Q$, and feedforward mapping matrices. Here we run Turnip and ZeRO Infinity (Rajbhandari et al., 2021); ZeRO is executed using all three “stages” (gradient partitioning, model weight partitioning, optimizer state partitioning) as well as CPU offload. Both Turnip and ZeRO use checkpointing during the forward pass to reduce the memory footprint. We measure the time it takes to run the forward and backward pass for one batch, with varying batch sizes and sequence lengths. All Turnip model weights are stored as single precision (32 bits). Results for training using both the P100 and the A100 server at 7B parameters are given in Figure 13. Both systems had a difficult time training the 65B parameter model. Turnip was faster for the one case it was able to run (1K length sequence, batch size eight took 58.5 seconds using Turnip and 72.9 using ZeRO Infinity) but Zero Infinity was more robust to larger batch sizes, where Turnip failed.
(3) The effect of nondeterminism. One of the key hypotheses of this paper is that nondeterminism can help performance, by allowing the system to react to the observed state of the computation. To test this, we implement a (semi-)deterministic version of Turnip and run the same set of experiments using this new version. Specifically, we add dependencies to fix the order of operations at each of the GPUs to be exactly the order that is input into BuildMemgraph. Unfortunately, this does not remove all nondeterminism from the system because of the way that reductions (summations) of tensors are handled by the system; fully removing all nondeterminism from Turnip is not easily possible. Still the experiment may be instructive. Figure 14 shows the speedup obtained by the nondeterministic version compared to the deterministic, over all experiments.
Discussion. Throughout the first token inference experiments, Turnip typically performed the best in terms of latency, with ZeRO Inference generally much slower but outperforming FlexGen, except for larger sequence sizes. This would seem to validate the highly non-deterministic, dataflow-based approach advocated for in the paper, at least if the goal is low latency.
To be fair, we note that FlexGen is designed for high throughput, as opposed to low latency, and FlexGen utilizes multiple GPUs only via pipelined parallelism. Note that FlexGen does not seem to get any slower when moving from batch size of eight to 16 on the A100 server. This suggests that filling the pipeline leads to substantial latency. Further, pipelined parallelism is more effective with more work in each pipeline stage, due to the high synchronization overhead and the need to try to overlap communication with computation, perhaps explaining FlexGen’s better performance for larger sequences, which are more dense computationally.
ZeRO Inference takes a much different approach, but it uses a highly synchronized form of model parallelism as it traverses the levels in a transformer, also carefully trying to overlap communication and computation, which may simply be more effective when there is more work to at each level. Turnip, on the other hand, is radically different. It does not “understand” the levels in a transformer, does not need to synchronize processing of the various levels, and simply tries to asynchronously process kernels as fast as it can. If it is stuck waiting for communication, it simply tries to do something else.
For training, there were clear advantages of Turnip over ZeRO Infinity, especially for smaller sequence lengths. This was particularly true on the A100 server, where Turnip was often much faster. For batches of sequences of length 1K, Turnip often took less than 50% of the time to process each batch, compared to ZeRO (at a batch size of 8, the time to process 1K sequences was 5.6 seconds for Turnip and 12.5 seconds for ZeRO, for batch size of 32 it was 12.1 seconds for Turnip and 17.8 seconds for ZeRO). The differences in performance were much less pronounced on the P100 server, though there Turnip was still faster. Finally, we note that both systems suffered significant out-of-memory errors during training. Interestingly, Turnip seemed to have more problems with memory on the A100 server, whereas ZeRO Infinity had more problems with memory on the P100 server. We conjecture that some of that could be solved in Turnip with a better input dataflow graph, which cuts the input problem into smaller pieces.
@@ 9 Conclusions
Appendix A Turnip Engine Details